OpenStudy (anonymous):

what are the critical points for F(x) = -2*(x^2+1)/(x^2-1)^2

OpenStudy (anonymous):

Agh. D: This one's messy. Get some quotient rule & chain rule in there, take the derivative and set it equal to zero. Keep the -2 on the outside the whole time, for simplicity.

myininaya (myininaya):

also find where f' doesnt exist

OpenStudy (anonymous):

dy^2/d^2x -2/(x^2-1) = dy/dx(4x)/(x^2-1) = (-12x^3-4)/(x^2 - 1)^3 set dy^2/d^2x to zero which equal

OpenStudy (anonymous):

Ah yes, and the second derivative as well. Points of inflection. :P

OpenStudy (anonymous):

wait oops.... i was look for the point of inflection sorry. So.....dy/dx -2/(x^2-1) = (4x)/(x^2-1) set the derivative to zero and solve. the first derivative is 0.

OpenStudy (anonymous):

there isn't a point of inflection because it is a monotonic function

OpenStudy (anonymous):

You're correct, my bad. :P Got me confused for a second there.

OpenStudy (anonymous):

complicated